# On the Importance of Embedding Norms in Self-Supervised Learning

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Self-supervised learning (SSL) has emerged as a powerful tool for learning representations from unlabeled data. This is because, when trained on large unlabeled datasets with contrastive objectives, SSL models often achieve performance levels comparable to those of supervised methods and can even induce emergent properties (Chen et al., 2020a; Caron et al., 2021). SSL has also significantly advanced multi-modal learning, particularly in vision-language models (Girdhar et al., 2023; Bommasani et al., 2021). Indeed, prominent methods such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and Florence (Yuan et al., 2021) all rely on standard SSL objectives discussed in this paper.
@@
Figure 1: Left: Network outputs (blue) lie in ambient $\mathbb{R}^{d}$, but only their projections (white) onto the hypersphere $S^{d-1}$ enter the loss. However, the embedding norms (shades of blue) influence the norms of the SSL loss gradient. Middle: Gradient field of the cosine similarity with respect to the orange direction. Highlighted gradients (black) illustrate the inverse relationship between the gradient and embedding norm (Proposition 1) The blue points trace an embedding’s trajectory using gradient descent. Right: Because gradient updates are orthogonal to an embedding point, they must increase the point’s norm (Corollary 3.3).
At the same time, there has been some empirical evidence that the embedding norms prior to normalization contain meaningful information. For example, Kirchhof et al. (2022) noted that embedding norms are related to network confidence. No theoretical explanation of this phenomenon exists. Separately, Zhang et al. (2020) and Wang et al. (2017) showed that training based on the cosine similarity loss affects the embedding norms and that the embedding norms affect the gradient magnitudes. However, the implications of this for SSL methods have not been fully explored. This implies a clear gap in the literature: several works suggest interactions between SSL models and the embedding norms but the extent of this relationship is unknown. Our work provides the first thorough analysis showing how the embedding norms interact with SSL training dynamics.
First, we prove theoretical bounds showing that the embedding norms impose a quadratic slowdown on SSL convergence and verify that this occurs in simulation and in practice. At the same time, we show that optimizing the cosine similarity grows these embedding norms across SSL methods. This creates an inherent catch-22: small embedding norms are required to train SSL models but these norms grow during training. Put simply, effectively training SSL models requires managing the embedding norms. We offer and evaluate several mechanisms for doing this.
Second, we argue that, because the embeddings grow with each gradient update, their norms naturally correspond to the frequency of observed latent features. Since models are more certain in frequently observed data (Zhong et al., 2021), we contend that SSL embedding norms encode a model’s confidence. We devise methods for studying this and show that it holds with remarkable consistency.
@@
Our code is available at https://github.com/Andrew-Draganov/SSLEmbeddingNorms. This paper supersedes our earlier workshop paper (Draganov et al., 2024).
@@ 2.1 Preliminaries & SSL methods
$\mathcal{L}_{ij}(\mathbf{Z})$ $=-\log\frac{\text{ExpSim}(z_{i},z_{j})}{\sum_{k\not\sim i}\text{ExpSim}(z_{i},z_{k})}$ (1)
$=-\underbrace{\hat{z}_{i}^{\top}\hat{z}_{j}}_{\mathcal{L}^{\mathcal{A}}_{ij}(\mathbf{Z})}+\underbrace{\log\left(S_{i}\right)}_{\mathcal{L}^{\mathcal{R}}_{ij}(\mathbf{Z})},$
where $\hat{z}$ is the unit vector of $z$, $\text{ExpSim}(a,b)=\exp\big{(}\hat{a}^{\top}\hat{b}\big{)}$ is the exponent of the cosine similarity, and $S_{i}=\sum_{k\not\sim i}\text{ExpSim}(z_{i},z_{k})$ represents the sum over all $k\neq i$ in the batch. For ease of interpretation, we split the loss term into attractive and repulsive components, resp. $\mathcal{L}^{\mathcal{A}}_{ij}$ and $\mathcal{L}^{\mathcal{R}}_{ij}(\mathbf{Z})$. Note that $\mathcal{L}^{\mathcal{A}}_{ij}(\mathbf{Z})$ is simply the negative cosine similarity between positive pair $z_{i}$ and $z_{j}$.
@@
Curiously, methods like BYOL (Grill et al., 2020) and SimSiam (Chen & He, 2021) showed that one can simply optimize $\mathcal{L}^{\mathcal{A}}_{ij}$, and avoid collapse by applying only the gradients to embedding $z_{i}$ (rather than to both $z_{i}$ and $z_{j}$). We refer to these as non-contrastive methods. Throughout this paper, we will use SimCLR with the InfoNCE loss to represent a prototypical contrastive SSL model and SimSiam with the negative cosine similarity to represent a prototypical non-contrastive SSL model. As is standard (Oquab et al., 2023), we use the $k$-nn classifier accuracy on the embedding space to evaluate the quality of a learned representation.
@@ 2.2 Related Work
Analysis of SSL Embeddings.
Due to its reliance on the cosine similarity, the seminal work of Wang & Isola (2020) showed that contrastive learning must satisfy two requirements: all positive pairs must be near one another (alignment) and all negative samples must spread evenly over the hypersphere (uniformity). Expanding on this blueprint, subsequent works have sought to formalize the learning capacity of contrastive methods (Saunshi et al., 2019; Wang & Isola, 2020; Zimmermann et al., 2021; HaoChen et al., 2021; Saunshi et al., 2022) while much of the research into non-contrastive methods has focused on how their architectures help to prevent collapse (Jing et al., 2021; Tian et al., 2021a; Zhang et al., 2022; Richemond et al., 2023).
SSL on Imbalanced Data.
@@
Relationship of Embedding Norms and SSL.
While SSL representation learning has largely focused on hyperspherical embedding distributions, some work has examined the embedding norm’s role. Wang et al. (2017); Zhang et al. (2018, 2020) all noted that embedding norms inversely scale gradients under the cosine similarity loss and suggested that the embeddings must grow. However, each paper largely brushed this interaction away: Wang et al. (2017) suggested that $\ell_{2}$-normalization suffices to handle the embedding norms, Zhang et al. (2018) states that attempts to resolve this interaction were unsuccessful and Zhang et al. (2020) provides a regularization term which shrinks the variance of the embedding norms but leaves their average magnitude unmanaged. This paper therefore differs significantly from the prior literature by (i) extending these results to the InfoNCE loss, (ii) evaluating the full extent to which large embedding norms affect real-world training, and (iii) providing principled mechanisms for addressing this effect.
Similarly, it has been suggested in Scott et al. (2021), Kirchhof et al. (2022) and Kirchhof et al. (2023b) that an SSL embedding’s magnitude could serve as a measure for the model’s certainty. This was evidenced in two ways: first, by qualitatively showing that samples with high-embedding norm are good representatives of the classes (see Figure S6) and, second, by finding that the embedding norm is smaller for out-of-distribution (OOD) samples. Importantly, no explanation was given for this phenomenon and, to our knowledge, it has not been systematically evaluated. Thus, our work expands the literature by comprehensively establishing that embedding norms encode SSL model confidence and providing a thorough analysis for the phenomenon.
@@ 3 The Properties of SSL Gradients
We begin by studying the gradients of the cosine similarity with respect to an arbitrary point $z_{i}$. Throughout this section, we refer to $\mathbf{Z}$ as any set of points in $\mathbb{R}^{d}$, with no other assumptions over the distribution. The gradient acting on one of these points has the following structure:
@@
$\nabla_{i}=\nabla_{i}^{\mathcal{A}}+\frac{1}{\|z_{i}\|}\cdot\sum_{k\not\sim i}\left(\hat{z}_{k}\cdot\frac{\text{\emph{ExpSim}}(z_{i},z_{k})}{S_{i}}\right)_{\perp z_{i}}.$ (2)
@@
This orthogonality has also been noted in Wang et al. (2017) and Zhang et al. (2018). As a direct consequence of this projection onto the tangent plane, applying the cosine similarity or InfoNCE gradients to a point must grow its magnitude (visualized in Figure 1, middle and right):
@@
Furthermore, Proposition 1 has direct implications for convergence rates under the cosine similarity. The gradient’s magnitude directly scales the learning rate, since $z_{i}^{\prime}=z_{i}+\gamma\nabla_{i}=z_{i}+\left(\gamma\cdot\|\nabla_{i}\|\right)\hat{\nabla}_{i}.$ Thus, Proposition 1 can be interpreted as saying that the embedding norm and the sin of the angle parameterize the model’s learning rate. Indeed, both quadratically slow down convergence:
@@
Put simply, the change in the cosine similarity via a step of gradient descent scales quadratically with the embedding’s norm and the sin of the angle to its positive counterpart.
@@ 4.2 Effect of SSL Training on the Embedding Norms
We now consider how SSL training affects the embedding norms via a simplified training setting where the data is generated from latent classes. Inspired by Zimmermann et al. (2021); Kirchhof et al. (2023a), consider an SSL dataset as a set of latent class distributions $\{\tilde{\mathcal{Z}}_{1},...,\tilde{\mathcal{Z}}_{k}\}$, where each $\tilde{\mathcal{Z}}_{i}$ is a probability distribution on the $d$-dimensional hypersphere $\mathbb{S}^{d}$. Let the observations $x\in\mathcal{X}\subset\mathbb{R}^{D}$ be obtained via a generating process $g:\mathbb{S}^{d}\rightarrow\mathbb{R}^{D}$. That is, our dataset is obtained by randomly choosing a probability distribution $\tilde{\mathcal{Z}}_{i}$, drawing a sample $\tilde{z}$ from it, and applying $g$ to $\tilde{z}$. Following Zimmermann et al. (2021); Kirchhof et al. (2023a), we are training a neural network $f:\mathcal{X}\to\mathbb{S}^{d}$ via contrastive learning to produce a learned latent embedding $f(\mathcal{X})$.
We analyze the relationship of parameterized SSL training to the embedding norms by simulating the above scenario. Specifically, we choose centers for 4 latent classes uniformly at random from $\mathbb{S}^{10}$. We then sample 4K points around these centers and normalize them to the hypersphere. From this, we produce the dataset via generating process $g:\mathbb{S}^{10}\subset\mathbb{R}^{11}\rightarrow\mathbb{R}^{64}$, where $g$ is given by multiplication by a random matrix. We finally train a 2-layer feedforward network with the supervised InfoNCE loss function [^2] The supervised InfoNCE loss explicitly chooses positive pairs as those which belong to the same class. on this dataset.
Figure 2(b) plots each embedding’s magnitude in the learned space as a function of (inverse) density in embedding space. We use the distance to an embedding’s $10^{\text{th}}$ nearest neighbor under the cosine similarity metric as a proxy for inverse density. We see that embeddings in dense regions of the embedding space tend to have higher norm. For instance, there are no embeddings which are both in a dense latent region (distance ~$0.1$ to the $10^{\text{th}}$ neighbor) and have small norm (~$6$). This follows from Corollary 3.3: dense regions of the embedding space receive the most gradient updates and, consequently, those embeddings will grow the most. We also modify the simulation for Figure 2(c) by providing a class imbalance parameter to the class distribution. Namely, class $1$ now has sample probability ~$1/2$, class $2$ has sample probability ~$1/4$, and so on. We then see that over the course of training, the mean embedding norms for frequent classes grow to higher values than those for sparse classes.
Takeaways.
@@ 5 Embedding Norm as Network Confidence
Embedding Norms Encode Novelty.
Figure 3 demonstrates how embedding norms characterize a sample’s “out-of-distributionness”. On the left side of the figure, we trained SimCLR and SimSiam models on the Cifar-10 train set for 512 epochs and evaluated them across different data splits, normalizing all embedding norms by the Cifar-10 train set mean. The results reveal a clear pattern: embedding norms decrease progressively with increasing distributional distance from the training data. For example, the Cifar-10 test set contains novel but distributionally similar samples and therefore results in only slightly reduced norms. In contrast, the Cifar-100 data splits exhibit substantially smaller norms due to their greater distributional shift. This relationship holds symmetrically when training on Cifar-100 and evaluating on Cifar-10, as seen on the right side of Figure 3.
Embedding Norms Encode Classification Accuracy.
Another measure of a network’s confidence in an embedding is the accuracy with which that sample is classified. To this end, we use the same experimental setup as above and train SimCLR and SimSiam on the Cifar-10, Cifar-100, and ImageNet100 datasets. [^3] We default to the ImageNet-100 split (Wang & Isola, 2020) at huggingface.co/datasets/clane9/imagenet-100. We then normalize the embedding magnitudes by the maximum across the dataset and bucket the embeddings into ranges of $0.05$, giving us 20 embedding buckets over the dataset. Figure 4 (left) then shows the per-bucket accuracy of a $k$-nn classifier which was fit on all the embeddings with respect to the cosine similarity metric. Indeed, we see that the $k$-nn classifier’s accuracy shows a clear monotonic trend with the embedding norms across datasets and SSL models.
Embedding Norms Encode Human Confidence.
Interestingly, not only does the embedding norm provide a measure for the sample’s novelty and its classification accuracy, but it also provides a signal for human labelers’ confidence and their agreement among one another. Using the Cifar-10-N and Cifar-100-N labels from Wei et al. (2021), where each training sample is labeled by the consensus label over multiple human annotators, Figure 4 (middle) shows higher embedding norms correspond to more accurate consensus labels. Similarly, the Cifar-10-H dataset from Peterson et al. (2019) provides $\approx 50$ human predictions for each image from the Cifar-10 test set, allowing us to evaluate the labelers’ entropy. Figure 4 (right) shows that, as the embedding norms grow, the human labels have less entropy and are therefore more likely to agree with one another.
Takeaways.

### v2

@@ Abstract
Self-Supervised Learning, Embedding Norms, Representation Learning
@@ 1 Introduction
Self-supervised learning (SSL) has emerged as a powerful tool for learning representations from unlabeled data. This is because, when trained on large unlabeled datasets with contrastive objectives, SSL models often achieve performance levels comparable to those of supervised methods and can even induce emergent properties (Chen et al., 2020a; Caron et al., 2021). SSL has also substantially advanced multi-modal learning, particularly in vision-language models (Girdhar et al., 2023; Bommasani et al., 2021). Indeed, prominent methods such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021) and Florence (Yuan et al., 2021) all rely on standard SSL objectives discussed in this paper.
@@
Figure 1: Left: Network outputs (blue) lie in ambient $\mathbb{R}^{d}$, but only their projections (white) onto the hypersphere $S^{d-1}$ enter the loss. However, the embedding norms (shades of blue) influence the norms of the SSL loss gradient. Middle: Gradient field of the cosine similarity with respect to the orange direction. Highlighted gradients (black) illustrate the inverse relationship between the gradient and embedding norm (Proposition 1). The blue points trace an embedding’s trajectory using gradient descent. Right: Because gradient updates are orthogonal to an embedding point, they must increase the point’s norm (Corollary 3.3).
At the same time, there has been some empirical evidence that the embedding norms prior to normalization contain meaningful information. For example, Kirchhof et al. (2022) noted that embedding norms are related to network confidence. No theoretical explanation of this phenomenon was given. Separately, Zhang et al. (2020) and Wang et al. (2017) showed that training based on the cosine similarity loss affects the embedding norms and that the embedding norms affect the gradient magnitudes. However, the implications of this for SSL methods have not been fully explored. This implies a clear gap in the literature: several works suggest interactions between SSL models and the embedding norms but the extent of this relationship is unknown. Our work provides the first thorough analysis showing how the embedding norms interact with SSL training dynamics.
First, we prove theoretical bounds showing that the embedding norms impose a quadratic slowdown on SSL convergence and verify that this occurs in simulation and in practice. At the same time, we show that optimizing the cosine similarity grows these embedding norms across SSL methods. This leads to a catch-22: small embedding norms are required to train SSL models but these norms grow during training. Put simply, effectively training SSL models requires managing the embedding norms. We offer and evaluate several mechanisms for doing this.
Second, we argue that, because the embeddings grow with each gradient update, their norms naturally correspond to the frequency of observed latent features. Since models are more certain in frequently observed data (Zhong et al., 2021), we contend that SSL embedding norms encode a model’s confidence. We show that it consistently holds across SSL methods.
@@
This paper supersedes our earlier workshop paper (Draganov et al., 2024). Our code is available at https://github.com/Andrew-Draganov/SSLEmbeddingNorms.
@@ 2.1 Preliminaries & SSL methods
$\mathcal{L}_{ij}(\mathbf{Z})$ $=-\log\frac{\text{ExpSim}(z_{i},z_{j})}{\sum_{k\neq i}\text{ExpSim}(z_{i},z_{k})}$ (1)
$=-\underbrace{\hat{z}_{i}^{\top}\hat{z}_{j}/\tau}_{\mathcal{L}^{\mathcal{A}}_{ij}(\mathbf{Z})}+\underbrace{\log\left(S_{i}\right)}_{\mathcal{L}^{\mathcal{R}}_{ij}(\mathbf{Z})},$
where $\hat{z}=z/\|z\|$ is the unit vector of $z$, $\text{ExpSim}(a,b)=\exp\big{(}\hat{a}^{\top}\hat{b}/\tau\big{)}$ is the exponent of the temperature-scaled cosine similarity, and $S_{i}=\sum_{k\neq i}\text{ExpSim}(z_{i},z_{k})$ represents the sum over all $k\neq i$ in the batch. For ease of interpretation, we split the loss term into attractive and repulsive components, resp. $\mathcal{L}^{\mathcal{A}}_{ij}(\mathbf{Z})$ and $\mathcal{L}^{\mathcal{R}}_{ij}(\mathbf{Z})$. Note that $\mathcal{L}^{\mathcal{A}}_{ij}(\mathbf{Z})$ is simply the negative cosine similarity between positive pair $z_{i}$ and $z_{j}$ scaled by $\tau$. We assume $\tau=1$ for simplicity; in SimCLR, $\tau=0.5$.
@@
Curiously, methods like BYOL (Grill et al., 2020) and SimSiam (Chen & He, 2021) showed that one can simply optimize $\mathcal{L}^{\mathcal{A}}_{ij}$, and avoid collapse by applying only the gradients to embedding $z_{i}$ (rather than to both $z_{i}$ and $z_{j}$). We refer to these as non-contrastive methods. Throughout this paper, we will use SimCLR with the InfoNCE loss to represent a prototypical contrastive SSL model and SimSiam with the negative cosine similarity to represent a prototypical non-contrastive SSL model. As is standard (Oquab et al., 2023), we use the $k$NN classifier accuracy on the embedding space to evaluate the quality of a learned representation. We provide linear probe results in Table S4 in the Appendix to show the generality of our results.
@@ 2.2 Related Work
Analysis of SSL Embeddings
Due to its reliance on the cosine similarity, the seminal work of Wang & Isola (2020) showed that contrastive learning must satisfy two requirements: all positive pairs must be near one another (alignment) and all negative samples must spread evenly over the hypersphere (uniformity). Expanding on this blueprint, subsequent works have sought to formalize the learning capacity of contrastive methods (Saunshi et al., 2019; Zimmermann et al., 2021; HaoChen et al., 2021; Saunshi et al., 2022) while much of the research into non-contrastive methods has focused on how their architectures help to prevent collapse (Jing et al., 2021; Tian et al., 2021a; Zhang et al., 2022; Richemond et al., 2023).
SSL on Imbalanced Data
@@
Relationship of Embedding Norms and SSL
While SSL representation learning has largely focused on hyperspherical embedding distributions, some work has examined the embedding norm’s role. Wang et al. (2017); Zhang et al. (2018, 2020) all noted that embedding norms inversely scale gradients under the cosine similarity loss and suggested that the embeddings must grow. However, each paper largely brushed this interaction away: Wang et al. (2017) suggested that $\ell_{2}$-normalization suffices to handle the embedding norms, Zhang et al. (2018) states that attempts to resolve this interaction were unsuccessful and Zhang et al. (2020) provides a regularization term which shrinks the variance of the embedding norms but leaves their average magnitude unmanaged. Our paper therefore differs from the prior literature by (i) extending these results to the InfoNCE loss, (ii) evaluating the full extent to which large embedding norms affect real-world training, and (iii) providing principled mechanisms for addressing this effect.
Similarly, it has been suggested (Scott et al., 2021; Kirchhof et al., 2022, 2023b) that an SSL embedding’s magnitude could serve as a measure for the model’s certainty. This was evidenced in two ways: first, by qualitatively showing that samples with high-embedding norm are good representatives of the classes (see Figure S6) and, second, by finding that the embedding norm is smaller for out-of-distribution (OOD) samples. Importantly, these references only show the existence of this relationship but, to our knowledge, there has not been an explanation provided for why this occurs nor a systematic evaluation of the settings in which norms encode model certainty. Thus, our work expands the literature by establishing how and why embedding norms encode SSL model confidence and providing a thorough empirical analysis for the phenomenon.
@@ 3 The Properties of SSL Gradients
We begin by studying the gradients of the cosine similarity with respect to an arbitrary point $z_{i}$. Throughout this section, we assume $\tau=1$ and refer to $\mathbf{Z}$ as any set of points in $\mathbb{R}^{d}$, with no other assumptions over the distribution. The gradient acting on one of these points has the following structure:
@@
$\nabla_{i}=\nabla_{i}^{\mathcal{A}}+\frac{1}{\|z_{i}\|}\cdot\sum_{k\neq i}\left(\hat{z}_{k}\cdot\frac{\text{\emph{ExpSim}}(z_{i},z_{k})}{S_{i}}\right)_{\perp z_{i}}.$ (2)
Let $z_{l}$, $l\neq i$, be a sample in the denominator of $\mathcal{L}_{ij}^{\mathcal{R}}(\mathbf{Z})$, then the gradient of $\mathcal{L}_{ij}(\mathbf{Z})$ with respect to $z_{l}$ is
$\nabla_{l}=-\frac{1}{\|z_{l}\|}\cdot\frac{\text{\emph{ExpSim}}(z_{i},z_{l})}{S_{i}}(\hat{z}_{i})_{\perp z_{l}}.$ (3)
@@
This orthogonality has also been noted in Wang et al. (2017) and Zhang et al. (2018).
As a direct consequence of this projection onto the tangent plane, applying the cosine similarity or InfoNCE gradients to a point must grow its magnitude (visualized in Figure 1, middle and right):
@@
Furthermore, Proposition 1 has direct implications for convergence rates under the cosine similarity. The gradient’s magnitude directly scales the learning rate, since $z_{i}^{\prime}=z_{i}+\gamma\nabla_{i}=z_{i}+\left(\gamma\cdot\|\nabla_{i}\|\right)\hat{\nabla}_{i}.$ Thus, Proposition 1 can be interpreted as saying that the embedding norm and the sine of the angle parameterize the model’s learning rate. Indeed, both quadratically slow down convergence:
@@
Put simply, the change in the cosine similarity via a step of gradient descent scales quadratically with the embedding’s norm and the sine of the angle to its positive counterpart.
@@ 4.1 Effect of the Embedding Norms on SSL Training
To further investigate the interaction between weight decay and the embedding-norm effect described in Theorem 3.4, we extend this simulation by augmenting the objective with a regularization term on the embedding norm. That is, we now optimize $\mathcal{L}_{ij}=-(z^{i})^{\top}z^{j}+\gamma\cdot\|z_{i}\|^{2}$, where $\gamma$ represents the weight decay strength. We initialize embedding pairs with different initial norms (1, 4, and 7) to systematically test the quadratic dependence predicted by our theory. The results, shown in Table 1, demonstrate that while weight decay can indeed accelerate convergence by controlling embedding growth, the fundamental quadratic dependence on initial embedding norm persists across all tested regularization strengths. For moderate weight decay values ($0.5\leq\gamma\leq 1$), we observe faster convergence compared to the unregularized case, with higher regularization providing greater speedup. However, when weight decay becomes too aggressive ($\gamma=10$), the representation collapses entirely, presumably because the embedding is more incentivized to shrink to the origin than to maximize the cosine similarity. These findings align with our practical experiments in Section 6, where excessive weight decay led to representation collapse.
Table 1: Steps to convergence for different initial embedding norms under varying weight decay strengths. The objective is modified to $\mathcal{L}_{ij}=-(z^{i})^{\top}z^{j}+\gamma\cdot\|z_{i}\|^{2}$ to simulate weight decay effects.
Initial Embedding Norm $\gamma=0.5$ (Steps) $\gamma=1.0$ (Steps) $\gamma=10.0$ (Steps)
1 64 49 divergence
4 526 318 divergence
7 1080 610 divergence
@@ 4.2 Effect of SSL Training on the Embedding Norms
We now consider how SSL training affects the embedding norms via a simplified training setting where the data is generated from latent classes. Inspired by Zimmermann et al. (2021) and Kirchhof et al. (2023a), consider an SSL dataset as a set of latent class distributions $\{\tilde{\mathcal{Z}}_{1},...,\tilde{\mathcal{Z}}_{k}\}$, where each $\tilde{\mathcal{Z}}_{i}$ is a probability distribution on the $d$-dimensional hypersphere $\mathbb{S}^{d}$. Let the observations $x\in\mathcal{X}\subset\mathbb{R}^{D}$ be obtained via a generating process $g:\mathbb{S}^{d}\rightarrow\mathbb{R}^{D}$. That is, our dataset is obtained by randomly choosing a distribution $\tilde{\mathcal{Z}}_{i}$, drawing a sample $\tilde{z}$ from it, and applying $g$ to $\tilde{z}$. We are training a neural network $f:\mathcal{X}\to\mathbb{S}^{d}$ via contrastive learning to produce a learned latent embedding $f(\mathcal{X})$.
We analyze the relationship of parameterized SSL training to the embedding norms by simulating the above scenario. Specifically, we choose centers for 4 latent classes uniformly at random from $\mathbb{S}^{10}$. We then sample 4K points around these centers and normalize them to the hypersphere. From this, we produce the dataset via generating process $g:\mathbb{S}^{10}\subset\mathbb{R}^{11}\rightarrow\mathbb{R}^{64}$, where $g$ is given by multiplication by a random matrix. We finally train a 2-layer feedforward network with the supervised InfoNCE loss function (which explicitly samples positive pairs belonging to the same class) on this dataset.
Figure 2(b) plots each embedding’s magnitude in the learned space as a function of (inverse) density in embedding space. We use the distance to an embedding’s $10^{\text{th}}$ nearest neighbor under the cosine similarity metric as a proxy for inverse density. We see that embeddings in dense regions of the embedding space tend to have higher norm. This follows from Corollary 3.3: dense regions of the embedding space receive the most gradient updates and, consequently, those embeddings will grow the most. We also modify the simulation for Figure 2(c) by providing a class imbalance parameter to the class distribution. Namely, class $1$ now has sample probability ~$1/2$, class $2$ has sample probability ~$1/4$, and so on. We then see that over the course of training, the mean embedding norms for frequent classes grow to higher values than those for sparse classes.
Takeaway
@@ 5 Embedding Norm as Network Confidence
Embedding Norms Encode Novelty
Figure 3 demonstrates how embedding norms characterize a sample’s “out-of-distributionness”. On the left side of the figure, we trained SimCLR and SimSiam models on the Cifar-10 train set for 512 epochs and then compared the embedding norms across different data splits, normalizing all embedding norms by the Cifar-10 train set mean. The results reveal a clear pattern: embedding norms decrease progressively with increasing distributional distance from the training data. For example, the Cifar-10 test set contains novel but distributionally similar samples and therefore results in only slightly reduced norms. In contrast, the Cifar-100 data splits exhibit substantially smaller norms due to their greater distributional shift. This relationship holds symmetrically when training on Cifar-100 and evaluating on Cifar-10, as seen on the right side of Figure 3.
Embedding Norms Encode Classification Accuracy
Another measure of a network’s confidence in an embedding is the accuracy with which that sample is classified. To this end, we use the same experimental setup as above and train SimCLR and SimSiam on the Cifar-10, Cifar-100, and ImageNet100 datasets. [^2] We default to the ImageNet-100 split (Wang & Isola, 2020) at huggingface.co/datasets/clane9/imagenet-100. We then normalize the embedding magnitudes by the maximum across the dataset and bucket the embeddings into ranges of $0.05$, giving us 20 embedding buckets over the dataset. Figure 4 (left) then shows the per-bucket accuracy of a $k$NN classifier which was fit on all the embeddings with respect to the cosine similarity metric. Indeed, we see that the $k$NN classifier’s accuracy shows a clear monotonic trend with the embedding norms across datasets and SSL models.
Embedding Norms Encode Human Confidence
Interestingly, not only does the embedding norm provide a measure for the sample’s novelty and its classification accuracy, but it also provides a signal for human labelers’ confidence and their agreement among one another. Using the Cifar-10-N and Cifar-100-N labels from Wei et al. (2021), where each training sample is labeled by the consensus label over multiple human annotators, Figure 4 (middle) shows higher embedding norms correspond to more accurate consensus labels. Similarly, the Cifar-10-H dataset from Peterson et al. (2019) provides ${\sim}50$ human predictions for each image from the Cifar-10 test set, allowing us to evaluate the entropy of the label distribution. Figure 4 (right) shows that, as the embedding norms grow, the human labels have less entropy (i.e., they are more likely to agree with one another).
Takeaways